# SplitFlux: Learning to Decouple Content and Style from a Single Image

### v1

@@ Abstract
Disentangling image content and style is essential for customized image generation. Existing SDXL-based methods struggle to achieve high-quality results, while the recently proposed Flux model fails to achieve effective content–style separation due to its underexplored characteristics. To address these challenges, we conduct a systematic analysis of Flux and make two key observations: (1) Single Dream Blocks are essential for image generation; and (2) Early single stream blocks mainly control content, whereas later blocks govern style. Based on these insights, we propose SplitFlux, which disentangles content and style by fine-tuning the single dream blocks via LoRA, enabling the disentangled content to be re-embedded into new contexts. It includes two key components: (1) Rank-Constrained Adaptation. To preserve content identity and structure, we compress the rank and amplify the magnitude of updates within specific blocks, preventing content leakage into style blocks. (2) Visual-Gated LoRA. We split the content LoRA into two branches with different ranks, guided by image saliency. The high-rank branch preserves primary subject information, while the low-rank branch encodes residual details, mitigating content overfitting and enabling seamless re-embedding. Extensive experiments demonstrate that SplitFlux consistently outperforms state-of-the-art methods, achieving superior content preservation and stylization quality across diverse scenarios.
@@ 1 Introduction
To address these limitations, we propose SplitFlux, a novel framework that disentangles content and style from a single image, enabling the disentangled content to be re-embedded into new contexts (see Fig. LABEL:fig:teaser). Inspired by B-LoRA [frenkel2024implicit], we perform extensive qualitative and quantitative analyzes of the Flux model to investigate the roles of different layers and derive two key findings: (1) Single Dream Blocks play a crucial role in image generation; (2) Early single stream blocks primarily control image content, whereas later blocks determine image style.
Based on these insights, we fine-tune the Single Dream Blocks of Flux via LoRA, which directly enables content–style disentanglement. However, this process may lead to the degradation of identity and structure in the disentangled content. To mitigate this, we introduce Rank-Constrained Adaptation that compresses the rank while amplifying update magnitude within specific blocks, preventing content leakage into style blocks and preserving identity and structure. Furthermore, to effectively re-embed the disentangled content into new contexts, we propose Visual-Gated LoRA. This module splits the content LoRA into two branches with different ranks and employs a routing function derived from image saliency to ensure that the high-rank branch captures the subject information, while the low-rank branch encodes residual details. The complementary loss encourages the two branches to learn diverse yet complementary representations, enabling the disentangled subject to be seamlessly integrated into new contexts. In summary, our key contributions are as follows:

### v2

@@ Abstract
Disentangling image content and style is essential for customized image generation. Existing SDXL-based methods struggle to achieve high-quality results, while the recently proposed Flux model fails to achieve effective content-style separation due to its underexplored characteristics. To address these challenges, we conduct a systematic analysis of Flux and make two key observations: (1) Single Stream Blocks are essential for image generation; and (2) Early single stream blocks mainly control content, whereas later blocks govern style. Based on these insights, we propose SplitFlux, which disentangles content and style by fine-tuning the single stream blocks via LoRA, enabling the disentangled content to be re-embedded into new contexts. It includes two key components: (1) Rank-Constrained Adaptation. To preserve content identity and structure, we compress the rank and amplify the magnitude of updates within specific blocks, preventing content leakage into style blocks. (2) Visual-Gated LoRA. We split the content LoRA into two branches with different ranks, guided by image saliency. The high-rank branch preserves primary subject information, while the low-rank branch encodes residual details, mitigating content overfitting and enabling seamless re-embedding. Extensive experiments demonstrate that SplitFlux consistently outperforms state-of-the-art methods, achieving superior content preservation and stylization quality across diverse scenarios.
@@ 1 Introduction
To address these limitations, we propose SplitFlux, a novel framework that disentangles content and style from a single image, enabling the disentangled content to be re-embedded into new contexts (see Fig. LABEL:fig:teaser). Inspired by B-LoRA [frenkel2024implicit], we perform extensive qualitative and quantitative analyzes of the Flux model to investigate the roles of different layers and derive two key findings: (1) Single Stream Blocks play a crucial role in image generation; (2) Early single stream blocks primarily control image content, whereas later blocks determine image style.
Based on these insights, we fine-tune the Single Stream Blocks of Flux via LoRA, which directly enables content–style disentanglement. However, this process may lead to the degradation of identity and structure in the disentangled content. To mitigate this, we introduce Rank-Constrained Adaptation that compresses the rank while amplifying update magnitude within specific blocks, preventing content leakage into style blocks and preserving identity and structure. Furthermore, to effectively re-embed the disentangled content into new contexts, we propose Visual-Gated LoRA. This module splits the content LoRA into two branches with different ranks and employs a routing function derived from image saliency to ensure that the high-rank branch captures the subject information, while the low-rank branch encodes residual details. The complementary loss encourages the two branches to learn diverse yet complementary representations, enabling the disentangled subject to be seamlessly integrated into new contexts. In summary, our key contributions are as follows: